# From Laboratory to Real World: A New Benchmark Towards Privacy-Preserved Visible-Infrared Person Re-Identification

### v1

@@ 3.2 Benchmark Protocols
Considering real-world conditions, we design three protocols, i.e., camera independence (CI), entity independence (EI) and entity sharing (ES), to simulate different privacy-restricted scenarios.
3.2.1 Protocol 1: Camera Independence (CI)
@@
CI: all camera data remains isolated, which simulates the most strict setting where no data sharing occurs across every single device.
It solves the concerns about data privacy, ownership regulations, or security policies in reality. However, existing VI-ReID methods cannot be directly applied under CI due to the modality incomplete issue. Specifically, current VI-ReID methods require paired visible and infrared images of the same identity as input during training, which is achieved through the PK strategy. $P$ identities with $K$ visible and $K$ infrared images are randomly sampled from the training set (i.e., mini-batch=$2\times P\times K$). These methods typically use a dual-stream architecture to extract modality-specific features in the shallow layers and modality-shared features in the deeper layers, relying on modality information as a prerequisite. Consequently, these requirements make existing VI-ReID methods incompatible with the CI protocol due to the challenges posed by modality incomplete.
@@
To solve these issues, we convert the two-stream structure into one-stream architecture and only sample images according to identity rather than identity and modality. Specifically, we randomly sample $K$ images from $P$ identities in the m-th local private data $D_{m}$, where the $K$ images can be visible, infrared, or a mixed of both. This eliminates the reliance on the paired visible and infrared images of the same subject as the input, thereby addressing the modality incomplete. In addition, we also employ channel augmentation (CA) [43] to destroy the color information by randomly selecting one channel (R, G, or B) to replace the other channels, which reduces the cross-modality gap across clients.
3.2.2 Protocol 2&3: Entity Independence and Entity Sharing (EI&ES)
@@
EI: Data from different entities is kept isolated during training, and the model is trained using a decentralized way. Then evaluate the trained model on an unseen entity to assess its generalization ability.
@@
In such an ES protocol, existing VI-ReID methods can be seamlessly applied. However, note that these methods’ evaluating protocols are often designed on the same dataset, where the data distribution is assumed to be consistent. Meanwhile, in our provided ES protocol, we merge several existing datasets and conduct the evaluation in a cross-domain manner to simulate real-world scenarios. To the best of our knowledge, this has not been explored and remains an open issue.
We reproduce several VI-ReID methods [26, 44, 43, 49, 14] under this protocol, which will be discussed in Sec. 4.3. Additionally, we provide a baseline using Empirical Risk Minimization (ERM) [36]. It merges data from all training domains to learn domain-invariant knowledge and can be formulated as follows:
@@
3.3 DPPT Method
To address the challenges in CI and EI, we propose decentralized privacy-preserved training (DPPT). DPPT consists of five steps aimed at addressing the challenges encountered in these protocols that cannot be effectively resolved through existing VI-ReID methods:
We first outline the challenges encountered under CI and EI, which cannot be handled by existing VI-ReID methods:
  • Domain shift: $P_{i}(x|y)\neq P_{j}(x|y)$. Pedestrian images on different clients will exhibit the non-iid (Independent and identically distributed) issue, where the corresponding feature distributions are different across different clients.
  • Identity missing: $P_{i}(y)\neq P_{j}(y)$. The subject is hard to appear in all cameras or entities.
Therefore, we propose the DPPT, which consists of five steps that we will introduce one by one.
Local Training: The process of client local training is shown in Fig. 2(b). The private data are fed into the encoder and get the $l_{2}$-normalized feature embedding $z$ and logits $l$. The identity loss $\mathcal{L}_{id}$ and circle loss $\mathcal{L}_{cir}$ are employed as the baseline loss functions (provided in the supplementary material) to ensure the learned features and logits are identity-related.
However, this training objective design cannot address the domain shift and identity missing, as the model tends to overfit the current private data in the client training and forget previously learned knowledge after download. Driven by this, we propose a memory rectification bank (MRB) that enables different clients to optimize towards the same goal.
@@ 3.3 DPPT Method
$\mathcal{L}_{mrb}=\frac{1}{N_{m}}\sum_{i=1}^{N_{m}}(1-\frac{\langle z_{i}^{m}\rangle^{(t)}\cdot\langle c_{y_{i}}^{g}\rangle^{(t-1)}}{||\langle z_{i}^{m}\rangle^{(t)}||\times||\langle c_{y_{i}}^{g}\rangle^{(t-1)}||}),$ (6)
where $\langle z_{i}^{m}\rangle^{(t)}$ is the i-th $l2$-normalized feature embedding in the t-th epoch on the m-th client. $\langle c_{y_{i}}^{g}\rangle^{(t-1)}$ is the global center of identity $y_{i}$ in the last epoch $t-1$, which contains the knowledge from the client $m$ and the other clients without leaking raw data. $\mathcal{L}_{mrb}$ pulls together the local memory and the global memory, addressing the domain shift and identity missing. The total training loss of the $m-th$ client can be written as:
@@
where $N_{total}$=$\sum_{i=1}^{n}N_{i}$ is the total number of samples across all clients.
@@
Input : The number of clients $n$, initial global model parameters $\theta^{g}$, private local model parameters $\{\theta^{l}_{1},\theta^{l}_{2},...,\theta^{l}_{n}\}$, local private datasets $\{D_{1},D_{2},...,D_{n}\}$, and the number of training epochs $E$.
Output : The global model parameters $\theta^{g}$ for test.
@@
       for $m=1,2,\dots,n$ do
             $\theta_{m}^{l},\mathcal{M}_{m}^{l}\leftarrow\textbf{ClientTraining($\theta^{g},D_{m},\mathcal{M}^{g}$)}$;
@@
       $\theta^{g}\leftarrow\sum_{i=1}^{n}\frac{|D_{i}|}{\sum_{j=1}^{n}|D_{j}|}\theta_{i}^{l}$ ;
@@
       $\mathcal{M}^{g}\leftarrow\{c_{i}^{g}=\frac{\sum_{m=1}^{n}\mathbbm{1}_{\{\hat{c}_{i}^{m}\neq 0\}}\cdot\hat{c}_{i}^{m}}{\sum_{m=1}^{n}\mathbbm{1}_{\{\hat{c}_{i}^{m}\neq 0\}}}|i=1,2,...,I\}$;
@@
$\theta_{m}^{l}\leftarrow\theta^{g}$ // Download the global model to the local model
@@
       $z_{i}=f_{m}(x_{i}),l_{i}=g_{m}(z_{i})$;
       $\mathcal{L}_{cir}\leftarrow(z_{i},y_{i})$, $\mathcal{L}_{id}\leftarrow(l_{i},y_{i})$ ;
       $\mathcal{M}_{m}^{l}\leftarrow z_{i}$ via Eq.(3,4);
       $\mathcal{L}_{mrb}\leftarrow(\mathcal{M}_{m}^{l},\mathcal{M}^{g})$ via Eq.(6);
       $\mathcal{L}=\mathcal{L}_{id}+\mathcal{L}_{cir}+\lambda\mathcal{L}_{mrb}$;
       $\theta^{l}_{m}\leftarrow\theta^{l}_{m}-\eta\nabla\mathcal{L}$;
return $\theta_{m}^{l},\mathcal{M}_{m}^{l}$;
Algorithm 1 DPPT
4 Experiment
@@
Table 1: Evaluation of our DPPT under CI protocol. Note that existing VI-ReID methods cannot directly be applied to the CI protocol as their frameworks are designed for data-shared learning. Thus, we implemented four classic federated learning algorithms as baselines to verify the efficacy of our method: FedProx [19], Fednova [37], Moon [18], and FedAvg [23]. Evaluating metrics rank-1(%), rank-10(%), mAP(%), and mINP(%) are reported. AGW and DNS are the reproduced VI-ReID method that removes the modality information.
@@
+DPPT (Ours) 51.27 88.55 49.29 34.47 59.85 77.58 55.58 41.70 34.69 67.20 41.91 38.48
Table 2: Evaluation of our DPPT under EI and ES protocols. The upper part of the table is under ES, while the lower part is under EI. The underlined and bold indicate the best results for both protocols, respectively. B is the baseline using ERM with $\mathcal{L}_{id}$ and $\mathcal{L}_{cir}$ under ES, B denote the baseline using FedAvg supervised by $\mathcal{L}_{id}$ and $\mathcal{L}_{cir}$ under EI. We use R, L, and S to denote RegDB, LLCM, and SYSU-MM01 datasets, respectively. The left of $\rightarrow$ indicates seen entities and the right is the unseen entity.
@@
LBA [26] 23.55 10.36 8.09 34.63 9.55 4.29 12.01 28.65 11.69 6.01 8.38 26.80 12.20 9.90
AGW [44] 23.55 10.36 9.59 38.28 10.42 4.45 13.79 29.76 14.05 8.25 9.08 26.77 12.37 9.80
DEEN [49] 41.23 27.70 9.48 38.39 10.03 3.72 $\underline{19.42}$ 37.59 $\underline{19.44}$ $\underline{12.33}$ 10.50 29.41 14.09 11.48
CAJ [43] 23.55 10.36 10.90 40.57 11.18 4.39 16.55 37.23 17.40 10.74 $\underline{11.35}$ $\underline{31.17}$ $\underline{15.03}$ $\underline{12.10}$
DNS [14] 25.45 10.36 $\underline{11.75}$ $\underline{42.36}$ $\underline{11.77}$ $\underline{4.56}$ 18.87 $\underline{37.79}$ 18.36 10.56 10.14 28.47 13.55 10.94
@@
B+CA 23.50 5.17 10.10 39.34 10.73 4.19 17.61 33.58 17.73 11.31 14.40 35.87 18.89 16.30
DPPT (Ours) 23.50 5.17 11.27 41.20 11.86 5.34 21.54 40.37 20.72 12.78 14.63 37.03 19.15 16.44
4.1 Experimental Settings
Datasets. We conduct experiments on three widely-used VI-ReID datasets to evaluate our proposed method, i.e., SYSU-MM01 [38], RegDB [24], and LLCM [49].
SYSU-MM01 contains 44,754 images of 491 identities captured by 4 visible and 2 infrared cameras, including 29,033 visible images and 15,715 infrared images. RegDB contains 412 identities collected by 1 visible and 1 infrared camera. Each identity has 10 visible and 10 infrared images. LLCM contains 46,767 images of 1,064 identities captured by 9 visible and infrared cameras, with the training and testing sets split at an approximate ratio of 2:1.
Evaluation Protocols. Following existing VI-ReID settings [49], we adopt the rank-k matching accuracy, mean Average Precision (mAP), and mean Inverse Negative Precision (mINP) [44] as evaluation metrics. The protocol in terms of CI is defined in Sec. 3.2.1, and the protocols of EI and ES are defined in Sec. 3.2.2. Following [44, 43, 49, 14, 38], all datasets are used in infrared-to-visible mode for clarity, where infrared images serve as the query set, and visible images as the gallery. All the reported results are repeated ten times for a fair comparison.
4.2 Implementation Details
@@
We adopt the ResNet-50 [8] pretrained on the ImageNet-1k [5] as our backbone. During the training state, all the input images are resized to $288\times 144$. The batch-size is set to 64, where randomly sampling 8 identities with 8 images. For the optimization, the Stochastic Gradient Descent (SGD) optimizer is adopted, with a weight decay of $5e^{-4}$ and momentum of 0.9. The initial learning rate is set to 0.2, and the OneCycleLR scheduler [32] is adopted. The epochs for CI and EC experiments are set to 50 and 30, respectively. The FedAvg [23] is employed as our default. The TOP $K=4$ and $\lambda=1$ are decided by ablations on settings.
@@
For CI, each client model is trained using data from a single camera within the training set of a given dataset. The global model is then evaluated on the testing set of the same dataset. For EI, we have two client models. Two datasets (entities) are used to train the two client models, respectively, and another new entity is used to test the global model. For ES, which is similar to EI, two entities are merged to train a model, and the model is tested on another new entity, which is not learned during training.
Figure 3: t-SNE and intra-inter distances visualization of each component. The circles and triangles denote the visible and infrared modality, respectively. Different colors denote different identities.
4.3 Qualitative Analysis

### v2

@@ III-B Benchmark Protocols
Considering real-world conditions, we design three protocols, i.e., camera independence (CI), entity independence (EI), and entity sharing (ES), to simulate different privacy-restricted scenarios.
@@
Protocol 1: Camera Independence (CI): all camera data remains isolated, which simulates the most strict setting where no data sharing occurs across every single device.
[…]
@@
To solve these issues, we convert the two-stream structure into a one-stream architecture and only sample images according to identity rather than identity and modality. Specifically, we randomly sample $K$ images from $P$ identities in the m-th local private data $D_{m}$, where the $K$ images can be visible, infrared, or a mix of both. This eliminates the reliance on the paired visible and infrared images of the same subject as the input, thereby addressing the modality incomplete. In addition, we also employ channel augmentation (CA) [14] by randomly selecting one channel (R, G, or B) to replace the other two channels, which reduces the cross-modality gap across clients.
Protocol 2&3: Entity Independence and Entity Sharing (EI&ES).
@@
EI: Data from different entities is kept isolated during training, and the model is trained using a decentralized way. Then the trained model is evaluated on an unseen entity to assess its generalization ability.
@@
In such an ES protocol, existing VI-ReID methods can be seamlessly applied. However, note that these methods’ evaluation protocols are often designed on the same dataset, where the data distribution is assumed to be consistent. Meanwhile, in our ES protocol, we merge several existing datasets and conduct the evaluation in a cross-domain manner to simulate real-world scenarios. To the best of our knowledge, this has not been explored before and remains an open issue.
We reproduce several VI-ReID methods [15, 13, 14, 3, 16] under this protocol, which will be discussed in Sec. IV-C. Additionally, we provide a baseline using Empirical Risk Minimization (ERM) [43]. It merges data from all training domains to learn domain-invariant knowledge and can be formulated as follows:
@@
III-C Image-based Decentralized Privacy-Preserved Training
To address the challenges in CI and EI, we propose the image-based decentralized privacy-preserved training (I-DPPT) for image data. I-DPPT consists of five steps aimed at addressing the challenges encountered in these protocols that cannot be effectively resolved through existing VI-ReID methods. We will introduce one by one.
Local Training: The process of client local training is shown in Fig. 3(b). The private data are fed into the encoder and get the $l_{2}$-normalized feature embedding $z$ and logits $l$. The identity loss $\mathcal{L}_{id}$ and circle loss $\mathcal{L}_{cir}$ are employed as the baseline loss functions (provided in the supplementary material) to ensure the learned features and logits are identity-related. However, this training objective design cannot address the domain shift and identity missing, as the model tends to overfit the current private data in the client training and forget previously learned knowledge after download. Driven by this, we propose a memory rectification bank (MRB) that enables different clients to optimize towards the same goal.
@@ III-C Image-based Decentralized Privacy-Preserved Training
[…]
where $\langle z_{i}^{m}\rangle^{(t)}$ is the i-th $l2$-normalized feature embedding in the t-th epoch on the m-th client. $\langle c_{y_{i}}^{g}\rangle^{(t-1)}$ is the global center of identity $y_{i}$ in the last epoch t1t-1, which contains the knowledge from the client $m$ and the other clients without leaking raw data. $\mathcal{L}_{mrb}$ pulls together the local memory and the global memory, addressing the domain shift and identity missing. The total training loss of the mthm-th client can be written as:
@@
where $N_{total}$$\sum_{i=1}^{n}N_{i}$ is the total number of samples across all clients.
@@
Input: Number of clients $n$; initial global model $\theta^{g}$; local models $\{\theta^{l}_{1},\dots,\theta^{l}_{n}\}$; datasets $\{D_{1},\dots,D_{n}\}$; number of epochs $E$
Output: Final global model $\theta^{g}$ for test
@@
 for $m=1$ to $n$ do
    $\theta_{m}^{l},\mathcal{M}_{m}^{l}\leftarrow$ ClientTraining($\theta^{g}$, $D_{m}$, $\mathcal{M}^{g}$)
   end for
@@
 $\theta^{g}\leftarrow\sum_{i=1}^{n}\frac{|D_{i}|}{\sum_{j=1}^{n}|D_{j}|}\theta_{i}^{l}$
@@
 $\mathcal{M}^{g}\leftarrow\{c_{i}^{g}=\frac{\sum_{m=1}^{n}\mathbbm{1}_{\{\hat{c}_{i}^{m}\neq 0\}}\cdot\hat{c}_{i}^{m}}{\sum_{m=1}^{n}\mathbbm{1}_{\{\hat{c}_{i}^{m}\neq 0\}}}\mid i=1,2,...,I\}$
end for
@@
$\theta_{m}^{l}\leftarrow\theta^{g}$ // Download the global model
@@
 $z_{i}=f_{m}(x_{i})$, $l_{i}=g_{m}(z_{i})$
 $\mathcal{L}_{cir}\leftarrow(z_{i},y_{i})$$\mathcal{L}_{id}\leftarrow(l_{i},y_{i})$
  Update $\mathcal{M}_{m}^{l}$ with $z_{i}$ via Eq.(3, 4)
 $\mathcal{L}_{mrb}\leftarrow(\mathcal{M}_{m}^{l},\mathcal{M}^{g})$ via Eq.(6)
 $\mathcal{L}=\mathcal{L}_{id}+\mathcal{L}_{cir}+\lambda\mathcal{L}_{mrb}$
 $\theta_{m}^{l}\leftarrow\theta_{m}^{l}-\eta\nabla\mathcal{L}$
end for
return $\theta_{m}^{l},\mathcal{M}_{m}^{l}$
Algorithm 1 I-DPPT
III-D Video-based Decentralized Privacy-Preserved Training
Unlike image-based VI-ReID, video-based VI-ReID takes as input a sequence of tracklets corresponding to each subject, which inherently contains temporal information. Within a tracklet, the same person may exhibit different poses over time, and different individuals often have distinct walking patterns and styles. As a result, video-based VI-ReID is more challenging than its image-based counterpart, as it requires addressing not only cross-modality and intra-modality variations but also learning temporal-invariant representations. Since our approach is the first to consider VI-ReID in a realistic, video-based setting, we aim to make minimal modifications to existing image-based methods to enable their seamless adaptation to video-based VI-ReID. Therefore, we only modify the local training stage, keeping other settings (including the proposed MRB) identical to image-based VI-ReID. We name it video-based decentralized privacy-preserved training (V-DPPT), and the details are illustrated in Fig. 4.
Figure 4: The local training of our proposed video-based decentralized privacy-preserved training (V-DPPT).
Specifically, let $D_{m}^{video}=\{(X_{i},y_{i})\}_{i=1}^{N_{m}}$ denote the m-th local video-level private data, where $X_{i}\in\mathbb{R}^{T\times H\times W\times C}$ and $y_{i}$ denote the pedestrian videos and corresponding label, and $N_{m}$ is the local data scale. The input is fed into the encoder to get the $l_{2}$-normalized feature embedding $z^{t}$. Then we average the $z^{t}$ along the time dimension to obtain the temporal-level representation, as defined below:
$z_{m}=\frac{1}{T}\sum_{t=1}^{T}z_{m}^{t},$ (9)
where T denotes the total number of images in the sequence. Then $z$ is input to the classifier and the proposed MRB. Finally, the identity loss $\mathcal{L}_{id}$, circle loss $\mathcal{L}_{cir}$, and the proposed memory rectification bank loss $\mathcal{L}_{MRB}$ are employed to optimize the whole training stage.
It should be noted that, unlike image-based VI-ReID, the fundamental unit of video-based VI-ReID is a pedestrian sequence rather than an individual image. Consequently, the mini-batch construction is different from image-based VI-ReID. Specifically, we randomly sample $K$ sequences from $P$ identities in the m-th local private data $D_{m}^{video}$. There is a total of $K\times P\times T$ images in each mini-batch, where $T$ is the total number of images in the sequence. The modality information of each sequence is unknown, but all images within a sequence have the same modality. Besides this, all the settings are identical to the image-based VI-ReID.
IV Experiments
IV-A Datasets and Evaluation Metrics
Datasets. We conduct experiments on five widely used VI-ReID datasets to evaluate our proposed method. Three of them are image datasets, i.e., SYSU-MM01 [2], RegDB [44], and LLCM [3]. The remaining two are video datasets, i.e., HITSZ-VCM [19] and BUPTCampus [21].
SYSU-MM01 [2] contains 44,754 images of 491 identities captured by four visible and two infrared cameras, including 29,033 visible images and 15,715 infrared images. RegDB [44] contains 412 identities collected by one visible and one infrared camera. Each identity has 10 visible and 10 infrared images. LLCM [3] contains 46,767 images of 1,064 identities captured by 9 visible and infrared cameras, with the training and testing sets split at an approximate ratio of 2:1. HITSZ-VCM [19] contain 927 identities captured by six cameras. Every 24 consecutive images in a video are regarded as a tracklet for a person during the same period. There are 251,452 visible images and 211,807 infrared images in total, which can be divided into 11,785 and 10,078 tracklets, respectively. 500 identities are used as the training set, and the remaining 427 identities are used as the testing set. BUPTCampus [21] contains 3,080 identities and 16,826 video tracklets, divided into 1,074 identities for primary learning, 930 for auxiliary learning, and 1,076 for testing. In our experiments, we exclude the auxiliary identities.
Evaluation Metrics. Following existing VI-ReID settings [3], we adopt the rank-k matching accuracy, mean Average Precision (mAP), and mean Inverse Negative Precision (mINP) [13] as evaluation metrics. The protocols in terms of CI, EI, and ES are defined in Sec. III-B. Following [13, 14, 3, 16, 2, 19, 21], all datasets are used in infrared-to-visible mode for clarity, where infrared images serve as the query set, and visible images as the gallery.
@@ IV-A Datasets and Evaluation Metrics
TABLE II: Evaluation of our I-DPPT under CI protocol (Best results in bold). Note that existing VI-ReID methods cannot be directly applied to the CI protocol as their frameworks are designed for data-shared learning. Thus, we implemented four classic federated learning algorithms as baselines to verify the efficacy of our method: FedProx [25], Fednova [45], Moon [26], and FedAvg [24]. Evaluating metrics rank-1(%), rank-10(%), mAP(%), and mINP(%) are reported. AGW and DNS are the reproduced VI-ReID method that removes the modality information.
@@
+I-DPPT (Ours) 51.27 88.55 49.29 34.47 59.85 77.58 55.58 41.70 34.69 67.20 41.91 38.48
TABLE III: Comparison with recent VI-ReID methods based on centralized training. All the compared methods are the centralized training method, but our I-DPPT adopts the decentralized training. The underlined indicate the best results for the existing centralized training VI-ReID methods. $\triangle$ denote the retrieval accuracy gap between our I-DPPT and the existing best centralized method.
@@
HOSNet [46] AAAI24 75.60 74.20 93.30 $\underline{89.20}$ 56.40 63.20
DSAF [47] TMM25 76.65 73.24 92.62 86.37 57.34 $\underline{64.27}$
DNS [16] ECCV24 77.27 74.35 93.49 88.10 $\underline{57.45}$ 64.11
FDNM [7] TIFS25 $\underline{77.80}$ $\underline{75.10}$ $\underline{94.00}$ 88.70 56.60 62.70
@@
$\triangle$ - 26.53 25.81 34.15 33.62 22.76 22.36
IV-B Implementation Details
@@
We adopt the ResNet-50 [48] pretrained on the ImageNet-1k [49] as our backbone. During the training state, all the input images are resized to $288\times 144$. For image-based VI-ReID, the batch size is set to 64, where eight identities with eight images are randomly sampled. For VCM, the batch size is set to 64, where similarly eight identities with eight sequences are randomly sampled. The sequence length $T$ is set to 6 following the official [19]. For BUPTCampus, the batch size is set to 16, where 16 identities with one sequence are randomly sampled. The sequence length $T$ is set to 10 in accordance with the official setting [21]. The Stochastic Gradient Descent (SGD) optimizer is adopted, with a weight decay of $5e^{-4}$ and momentum of 0.9. The initial learning rate is set to 0.2, and the OneCycleLR scheduler [50] is adopted. The number of epochs for CI and EI experiments is set to 50 and 30 for image-level settings, respectively, and 100 for video-level settings. The FedAvg [24] is employed as our default. The TOP $K=4$ and $\lambda=1$ are decided by ablations on settings.
@@
For CI, each client model is trained using data from a single camera within the training set of a given dataset. The global model is then evaluated on the testing set of the same dataset. For EI under image-level VI-ReID, we have two client models. Two datasets (entities) are used to train the two client models, respectively, and another new entity is used to test the global model. For ES, which is similar to EI, two entities are merged to train a model, and the model is tested on a new entity, which is not learned during training. Due to the limited video-based VI-ReID datasets, the EI protocol cannot be directly implemented. Instead, we use the model weights trained under the CI protocol for direct testing on other unseen datasets. For the video-based VI-ReID ES protocol, we reproduce several video-based VI-ReID methods by training them on one dataset and directly evaluating them on another unseen dataset.
TABLE IV: Evaluation of our I-DPPT under EI and ES protocols. The upper part of the table is under ES, while the lower part is under EI. The underlined and bold indicate the best results for both protocols, respectively. B is the baseline using ERM with $\mathcal{L}_{id}$ and $\mathcal{L}_{cir}$ under ES, B denote the baseline using FedAvg supervised by $\mathcal{L}_{id}$ and $\mathcal{L}_{cir}$ under EI. We use R, L, and S to denote RegDB, LLCM, and SYSU-MM01 datasets, respectively. The left of $\rightarrow$ indicates seen entities, and the right is the unseen entity.
[…]
LBA [15] ES 10.36 8.09 34.63 9.55 4.29 12.01 28.65 11.69 6.01 8.38 26.80 12.20 9.90
AGW [13] ES 10.36 9.59 38.28 10.42 4.45 13.79 29.76 14.05 8.25 9.08 26.77 12.37 9.80
DEEN [3] ES 27.70 9.48 38.39 10.03 3.72 $\underline{19.42}$ 37.59 $\underline{19.44}$ $\underline{12.33}$ 10.50 29.41 14.09 11.48
CAJ [14] ES 10.36 10.90 40.57 11.18 4.39 16.55 37.23 17.40 10.74 $\underline{11.35}$ $\underline{31.17}$ $\underline{15.03}$ $\underline{12.10}$
DNS [16] ES 10.36 $\underline{11.75}$ $\underline{42.36}$ $\underline{11.77}$ $\underline{4.56}$ 18.87 $\underline{37.79}$ 18.36 10.56 10.14 28.47 13.55 10.94
[…]
B+CA EI 5.17 10.10 39.34 10.73 4.19 17.61 33.58 17.73 11.31 14.40 35.87 18.89 16.30
I-DPPT (Ours) EI 5.17 11.27 41.20 11.86 5.34 21.54 40.37 20.72 12.78 14.63 37.03 19.15 16.44
IV-C Quantitative Analysis for Image-based VI-ReID